OpenStudy (anonymous):

change from rectangular to cylindrical coordinates (1, -1, 4)

OpenStudy (stacey):

r^2 = x^2 + y^2 \[\theta=\tan^{-1} (y/x)\]

OpenStudy (anonymous):

\[(r, \theta, z ) = ( \sqrt{2} , - \frac{\pi}{4} , 4 ) \]